OpenStudy (anonymous):

how do you simplify 2x^2+3x-3? also, x^2+2x+1?

OpenStudy (bmorse):

nevermind let me redo it

OpenStudy (bmorse):

how is your teacher solving? using the quadratic equation?

OpenStudy (anonymous):

yeah, its part of a "quadratic inequalities" worksheet

OpenStudy (anonymous):

but so far, she has just been factoring it not using the quad formula ex: x^2+2x-8=0 so (x+4)(x-2)=0 so x can be -4 or 2

OpenStudy (anonymous):

but I guess you could do it both ways

OpenStudy (bmorse):

so 2x^2=a 3x=b -3=c =\[-3\pm \sqrt{(3)^2-4(2)(-3)}\div(2\times2)\]

OpenStudy (anonymous):

yeah but that's rad 33, I think its supposed to be nice #

OpenStudy (bmorse):

first you try -3+sqrt(9+24)/4 which is -1.56 and then -3-sqrt(9+24)/4 which is -4.436 so x=-1.56 and -4.436

OpenStudy (anonymous):

ok thx